# Understanding Co-speech Gestures in-the-wild

### v1

@@ Abstract
Co-speech gestures play a vital role in non-verbal communication. In this paper, we introduce a new framework for co-speech gesture understanding in the wild. Specifically, we propose three new tasks and benchmarks to evaluate a model’s capability to comprehend gesture-text-speech associations: (i) gesture-based retrieval, (ii) gestured word spotting, and (iii) active speaker detection using gestures. We present a new approach that learns a tri-modal speech-text-video-gesture representation to solve these tasks. By leveraging a combination of global phrase contrastive loss and local gesture-word coupling loss, we demonstrate that a strong gesture representation can be learned in a weakly supervised manner from videos in the wild. Our learned representations outperform previous methods, including large vision-language models (VLMs), across all three tasks. Further analysis reveals that speech and text modalities capture distinct gesture-related signals, underscoring the advantages of learning a shared tri-modal embedding space. The dataset, model, and code are available at: https://www.robots.ox.ac.uk/~vgg/research/jegal.
@@ 1 Introduction
In this paper, our objective is to learn and evaluate co-speech gesture representations. To this end, we propose three tasks and evaluation benchmarks that act as a proxy for assessing real world applications: (1) gesture-based cross-modal retrieval, (2) gesture word spotting, and (3) active speaker detection via gestures. We perform large-scale training featuring $\approx 7000$ speakers and evaluate on in-the-wild videos from the AVSpeech dataset [20].
Figure 1: Co-speech gestures supplement the spoken language – we show examples for three phrases here, with common words. Learning to associate gestures with the uttered phrases is essential for a holistic understanding of human communication.
@@
We make three key design choices in our approach that result in a strong tri-modal gesture representation. To start with, we learn gesture video representations from large-scale weak cross-modal supervision. The supervision is weak because we only use phrase-level speech audio and transcripts – since we do not have any information on which words in the speech are gestured for videos collected ‘in-the-wild’. Second, we obtain cross-modal supervision in the form of both audio and the corresponding text transcript (in Section 6.1, we demonstrate that speech and text modalities capture complementary gesture-related signals). Third, we introduce a new gesture-word alignment and spotting loss that strongly encourages learning of word-level correspondences.
To summarize, we make the following contributions: (i) we propose a new framework for co-speech gesture understanding with three new tasks and evaluation benchmarks; (ii) we learn a joint tri-modal embedding space in a weakly-supervised manner with a combination of global phrase-level objective, and a local word-level gesture coupling loss; (iii) we demonstrate that the learned JEGAL representation outperforms previous vision-language foundation models on the three gesture-centric tasks and is useful for practical applications.
@@ 2 Related Work
Spectrum of Human Gestures. Gestures can be classified broadly into four major classes [34]. Emblematics convey a clear symbolic meaning, e.g. a “thumbs-up”. Iconic gestures are used to convey meaning co-occurring with the articulated speech, e.g. “revolving door” is accompanied with the hands moving in a circular motion. Deictic gestures are pointing gestures using the index finger. The most common are the “beat” gestures which are co-speech gestures that are temporally aligned with the prosodic characters of human speech. They are prominent on lexically stressed syllables. Past works [55, 26] have studied how these gesture classes relate to the speech. Several papers have been attempted to recognize hand, head, and facial emblematic gestures [22, 18]. Recognizing deictic gestures can help in identifying the referring object being pointed to in a conversation, an essential part of human-robot interaction [33]. The other two gesture classes, i.e.,  beat and iconic, are quite diverse and are difficult to associate with a well-defined set of words or hand movements. This work takes a data-driven approach to learn gesture representations with the help of speech and natural language.
@@ 3 Method
Our goal is to learn co-speech gesture representations from speech and text supervision. Given a dataset G, S, L of gesture clips, the accompanying speech segments and their corresponding transcriptions, our goal is to learn gesture representations that capture the rich semantics (from text, usually) and utterance style (usually from speech) of what is being spoken.
@@ 3.7 Training Objective
$\mathcal{L}_{seq}=\frac{1}{N}\sum_{i=1}^{N}\left(\log\frac{\exp(\gamma\cdot cos(g_{i},c_{i}))}{\sum_{j=1}^{n}\exp(\gamma\cdot cos(g_{i},c_{j}))}\right)$ (2)
Local Gesture-word Coupling Loss. Pooling the word-level representations to compute the global phrase loss can lead to weak gesture-word associations (Tab 6, row $1$). However, directly training to match gestures to words is not possible – very few words are gestured in a given phrase, and we do not know which of them are. Thus, we devise a new strategy to learn word-level correspondences using phrase-level supervision. Given a pair of word-level gesture and speech-text ($\mathbf{g^{w}},\mathbf{c^{w}}$) embeddings, we first find the closest gesture $g^{w_{j}}$ for each speech-text word $c^{w_{i}}$. Our hypothesis is that a matching ($\mathbf{g^{w}},\mathbf{c^{w}}$) will have a higher number of strong word couplings than a non-matching pair. With this idea, we define the following scoring function and the gesture-word coupling loss:
@@
$\mathcal{L}_{couple}=\frac{1}{N}\sum_{i=1}^{N}\left(\log\frac{\exp(\gamma\cdot\lambda(g^{w}_{i},c^{w}_{i}))}{\sum_{j=1}^{N}\exp(\gamma\cdot\lambda(g^{w}_{i},c^{w}_{j}))}\right)$ (4)
@@ 3.8 Implementation Details
Training data. We train our model, JEGAL, on triplets of gesture clips, speech segments, and text transcriptions. For the gesture frame inputs, we resize the frames to $270\times 480$ pixels. We extract melspectrograms with a hop length of $10$ms. The word-aligned text transcriptions are tokenized into wordpiece tokens. Using the start-end time of the word boundaries, we randomly sample a video clip between $2-10$ seconds in length.
@@ 3.9 Training Datasets
We train our model using the following datasets: (i) PATS [24], (ii) LRS3 [4], and (iii) LRS3-Lang [5, 25]. By merging these three datasets, we compile a comprehensive multilingual gesture dataset suitable for both training and evaluation. The specifics of each dataset are outlined in Table 1. PATS [24] is a publicly available dataset comprising $250$ hours of video content from $25$ speakers sourced from diverse platforms such as lectures, talk-shows, YouTube, and televangelists. The LRS3 [4] dataset is composed of $400$ hours of TED talks spanning a broad spectrum of speakers and a rich vocabulary. We also use a subset of the LRS3-lang [5] dataset, containing $366$ hours across $4$ distinct languages. We acquired this dataset directly from the authors, as it has not been made publicly available yet.
@@
Table 1: We train and evaluate on multiple datasets consisting of $720$ hours of gesture clips spanning $5$ languages and $7000+$ speakers. For evaluation, we curate task-specific benchmarks from the publicly available AVSpeech [20] dataset.
Dataset split ##\## hours ##\## spk. avg. clip ##\## videos
@@
AVS-Ret test 0.26 420 1.85 500
AVS-Spot test 0.38 391 2.73 500
AVS-Asd test 0.53 500 3.80 500
@@ 4.1 Cross-modal Retrieval
4.2 Spotting Gestured Words
@@ 4.3 Active Speaker Detection
Method Mod. Speech-text to Gesture retrieval Gesture to Speech-text retrieval T A R@5 $\uparrow$ R@10$\uparrow$ R@25$\uparrow$ R@50$\uparrow$ MR $\downarrow$ R@5 $\uparrow$ R@10$\uparrow$ R@25$\uparrow$ R@50$\uparrow$ MR $\downarrow$ Random 1.00 2.00 5.00 10.00 250 1.00 2.00 5.00 10.00 250 Zero-shot Clip4Clip [31] 1.90 2.70 5.90 15.10 194 2.00 3.10 7.20 12.80 205 Language-Bind [65] 1.70 2.30 6.00 14.00 207 2.00 3.20 6.60 12.20 212 GestSync [25] 1.80 2.60 6.60 14.40 200 1.80 3.60 11.00 20.20 159 Fine-tuned Clip4Clip [31] 2.40 4.60 8.90 14.00 214 2.40 5.00 9.20 15.60 219 Language-Bind [65] 2.40 4.40 9.00 15.60 187 2.80 4.60 9.00 14.40 199 GestSync [25] 3.60 7.20 18.00 30.20 113 6.20 9.20 21.00 31.20 109 GestureDiffuClip [10] 8.90 11.67 22.80 33.20 96 9.70 14.30 24.10 36.00 94 Ours JEGAL 15.80 30.00 50.20 58.80 25 10.80 19.20 34.60 47.80 55 JEGAL 14.20 26.00 49.40 58.80 26 11.60 20.20 33.20 45.00 62 JEGAL 17.40 28.40 53.00 69.80 22 16.40 26.40 43.60 55.20 38
@@ 4.4 AVSpeech Test Benchmarks
AVS-Spot. To quantitatively evaluate the gesture spotting task, we manually curate a new test dataset where we search and annotate clips that clearly contain a word that is gestured. We obtain $500$ such clips, each containing a target word that is clearly gestured. The manual annotation process removes all kinds of label noise in the test set, allowing for a faithful evaluation of our newly defined gesture spotting task. Additionally, we also manually annotate these target words with binary “stress/emphasis” labels, which can have important cues about the gesture (Table 5).
AVS-Asd. To build the evaluation dataset for active speaker detection, we first choose $500$ “target” clips. For these target clips, we create three evaluation subsets, where we choose $P-1$ clips from different speakers, where $P=2,4,6$. We report the accuracy of detecting the correct target speaker out of the $P$ different speakers.
@@ 5.1 Baselines
For baselines, we report performance of zero-shot pre-trained video-language models [65, 58] and pre-trained GestSync [25], which learns gesture-audio correspondences by solving for audio-visual synchronization. We also report scores after fine-tuning all these models further on our training data for a fair comparison. In addition, we compare with the semantic encoder of GestureDiffuCLIP [10], by training it on our dataset.
@@ 5.2 Cross-modal Retrieval
In Table 2, we compare the performance of JEGAL against other baselines on the cross-modal retrieval task. Zero-shot evaluation of foundational vision-language models like LanguageBind [65] and Clip4Clip [31] leads to slightly higher than chance performance. These models are designed to capture different kinds of features: they cannot handle a large number of frames, and learn non-gesture attributes like identity and scene. Fine-tuning these models nearly doubles their performance on the task, but it is still far from the performance of JEGAL . GestSync [25] is trained with a gesture-related task, and hence performs nearly twice as well as the foundational vision-language models. However, since this network is trained to detect synchronization offsets in speech and video, its representations perform poorly for global semantic tasks like retrieval. This is also partly true for our model when we turn off the global phrase loss (Table 6 row $2$ vs row $3$). GestureDiffuCLIP’s semantic encoder [10] performs the best among the baselines, however, the lack of local word-level semantic supervision leads to an inferior performance compared to JEGAL .
Furthermore, none of the baseline approaches ingest and fuse multi-modal speech-text inputs. Our JEGAL model outperforms previous methods by a large margin. JEGAL can retrieve gestures from speech or text queries with similar performance. The opposite direction is also true, i.e. retrieving speech or text for a query gesture clip. Finally, retrieving with the fused speech-text representation, $c$ is clearly better than the unimodal variants, showing that the speech and text embeddings each encode information that is not present in the other modality.
@@
Clip4Clip [31] 7.62
Language-Bind [65] 9.11
GestSync [25] 5.12
@@
Clip4Clip [31] 14.79
Language-Bind [65] 15.68
GestSync [25] 10.41
GestureDiffuClip [10] 21.50
@@
JEGAL 60.40
JEGAL 40.40
JEGAL 63.00
5.3 Spotting Gestured Words
@@ 5.4 Active Speaker Detection
In Table 4, we show the accuracy of identifying the target speaker for a given text and/or speech segment. This task is different from our other tasks – it does not need a strong holistic understanding of the gesture sequence like retrieval, nor does it need semantic word-level understanding. It can simply be solved by checking for frame-level synchronization, which is exactly why GestSync [25], LanguageBind [65] baselines perform well in this task. Also, compared to the previous tasks, speech information is more useful here compared to text. While the text-only variant of JEGAL lags behind GestSync [25], the speech and speech-text variants actually emerges as the best performing approaches. This task shows that JEGAL learns gesture features that are useful for a wide variety of tasks, across both speech and text modalities.
@@
Clip4Clip [31] 56.4 31.2 23.0
Language-Bind [65] 60.1 38.9 27.2
GestSync [25] 57.3 35.1 26.2
@@
Clip4Clip [31] 60.0 40.1 28.8
Language-Bind [65] 62.5 41.4 28.2
GestSync [25] 72.8 57.2 48.0
GestureDiffuClip [10] 66.8 43.9 34.8
@@
JEGAL 69.4 48.6 37.4
JEGAL 77.2 60.0 52.2
JEGAL 80.8 66.4 59.0
@@ 6.1 Speech v/s Text Modalities
We have already seen how our model can flexibly leverage speech or text modality to solve three different kinds of tasks. In Fig 4, we show an example where audio-based gesture word spotting is successful, but text-based spotting is not. And we see that the speech clearly has the word “action” being emphasized. This example leads us to do a deeper analysis of gesture word spotting based on stress cues. We divide the binary stress labels to split the AVS-Spot test set ($500$ samples) into two subsets, one containing only samples with stressed/emphasized words ($100$ samples) and the other subset containing the remaining words ($400$ samples). In Table 5, we report the same spotting accuracy metric on these subsets separately. While we saw previously in Table 3 that speech-based gesture spotting is clearly inferior compared to text-based gesture spotting, it is not always the case. We see that the difference in spotting accuracy between the stressed and non-stressed words is higher for the speech modality. This indicates that speech modality pays more attention to emphasis of the word compared to the text modality.
Table 5: Stressed words are more likely to be spotted with speech-based spotting than non-stressed words. As seen in the column “$\Delta$”, the difference between stressed vs. non-stressed word spotting is higher for the speech modality.

### v2

@@ Abstract
Co-speech gestures play a vital role in non-verbal communication. In this paper, we introduce a new framework for co-speech gesture understanding in the wild. Specifically, we propose three new tasks and benchmarks to evaluate a model’s capability to comprehend gesture-speech-text associations: (i) gesture based retrieval, (ii) gesture word spotting, and (iii) active speaker detection using gestures. We present a new approach that learns a tri-modal video-gesture-speech-text representation to solve these tasks. By leveraging a combination of global phrase contrastive loss and local gesture-word coupling loss, we demonstrate that a strong gesture representation can be learned in a weakly supervised manner from videos in the wild. Our learned representations outperform previous methods, including large vision-language models (VLMs). Further analysis reveals that speech and text modalities capture distinct gesture related signals, underscoring the advantages of learning a shared tri-modal embedding space. The dataset, model, and code are available at: https://www.robots.ox.ac.uk/~vgg/research/jegal.
@@ 1 Introduction
In this paper, our objective is to learn and evaluate co-speech gesture representations. To this end, we propose three tasks and evaluation benchmarks that act as a proxy for assessing real world applications: (1) gesture based cross-modal retrieval, (2) gesture word spotting, and (3) active speaker detection via gestures. We perform large-scale training featuring $\approx 7000$ speakers and evaluate on in-the-wild videos from the AVSpeech dataset [18].
Figure 1: Co-speech gestures supplement the spoken language – we show examples for six phrases here, with common words. Learning to associate gestures with the uttered phrases is essential for a holistic understanding of human communication.
@@
We make three key design choices in our approach that result in a strong tri-modal gesture representation. To start with, we learn gesture video representations from large-scale weak cross-modal supervision. The supervision is weak because we only use phrase-level speech audio and transcripts – since we do not have any information on which words in the speech are gestured for videos collected ‘in-the-wild’. Second, we obtain cross-modal supervision in the form of both audio and the corresponding text transcript (in Section 6.1, we demonstrate that speech and text modalities capture complementary gesture-related signals). Third, we introduce a new gesture-word alignment and spotting loss that explicitly encourages learning of word-level correspondences.
To summarize, we make the following contributions: (i) we propose a new framework for co-speech gesture understanding with three new tasks and evaluation benchmarks; (ii) we learn a joint tri-modal embedding space in a weakly-supervised manner with a combination of global phrase-level objective, and a local word-level gesture coupling loss; (iii) we demonstrate that the learned JEGAL representation performs on par with vision-language foundation models on three gesture-centric tasks and is useful for practical applications.
@@ 2 Related Work
Spectrum of Human Gestures. Gestures can be classified broadly into four major classes [32]. Emblematics convey a clear symbolic meaning, e.g. a “thumbs-up”. Iconic gestures are used to convey meaning co-occurring with the articulated speech, e.g. “revolving door” is accompanied with the hands moving in a circular motion. Deictic gestures are pointing gestures using the index finger. The most common are the “beat” gestures which are co-speech gestures that are temporally aligned with the prosodic characters of human speech. They are prominent on lexically stressed syllables. Past works [54, 24] have studied how these gesture classes relate to the speech. Several works have attempted to recognize hand, head, and facial emblematic gestures [20, 16]. Recognizing deictic gestures can help in identifying the referring object being pointed to in a conversation, an essential part of human-robot interaction [31]. The other two gesture classes, i.e.,  beat and iconic, are quite diverse and are difficult to associate with a well-defined set of words or hand movements. This work takes a data-driven approach to learn gesture representations with the help of speech and natural language.
@@ 3 Method
Our goal is to learn co-speech gesture representations from speech and text supervision. Given a dataset G, S, L of gesture clips, the accompanying speech segments and their corresponding transcriptions, our goal is to learn gesture representations that capture the rich semantics (from text) and utterance style (from speech) of what is being spoken.
@@ 3.7 Training Objective
$\mathcal{L}_{seq}=-\frac{1}{N}\sum_{i=1}^{N}\left(\log\frac{\exp(\gamma\cdot cos(g_{i},c_{i}))}{\sum_{j=1}^{n}\exp(\gamma\cdot cos(g_{i},c_{j}))}\right)$ (2)
where $\gamma$ is the temperature and coscos is the cosine similarity.
Local Gesture-word Coupling Loss. Pooling the word-level representations to compute the global phrase loss can lead to weak gesture-word associations (Table 6, row $1$). However, directly training to match gestures to words is not possible – very few words are gestured in a given phrase, and we do not know which of them are. Thus, we devise a new strategy to learn word-level correspondences using phrase-level supervision. Given a pair of word-level gesture and speech-text ($\mathbf{g^{w}},\mathbf{c^{w}}$) embeddings, we first find the closest gesture $g^{w_{j}}$ for each speech-text word $c^{w_{i}}$. Our hypothesis is that a matching ($\mathbf{g^{w}},\mathbf{c^{w}}$) will have a higher number of strong word couplings than a non-matching pair. With this idea, we define the following scoring function and the gesture-word coupling loss:
@@
$\mathcal{L}_{couple}=-\frac{1}{N}\sum_{i=1}^{N}\left(\log\frac{\exp(\gamma\cdot\lambda(g^{w}_{i},c^{w}_{i}))}{\sum_{j=1}^{N}\exp(\gamma\cdot\lambda(g^{w}_{i},c^{w}_{j}))}\right)$ (4)
@@ 3.8 Implementation Details
Training data. We train our model, JEGAL, on triplets of gesture clips, speech segments, and text transcriptions. For the gesture frame inputs, we resize the frames to $270\times 480$ pixels. We extract melspectrograms with a hop length of 1010ms. The word-aligned text transcriptions are tokenized into wordpiece tokens. Using the start-end time of the word boundaries, we randomly sample a video clip between 2102-10 seconds in length.
@@ 3.9 Training Datasets
We train our model using the following datasets: (i) PATS [22], and (ii) a subset of the MultiVSR dataset [40]. The dataset specifics are outlined in Table 1. PATS [22] is a publicly available video dataset from $25$ speakers sourced from diverse platforms such as lectures, talk-shows, YouTube, and televangelists. The subset from the MultiVSR dataset is composed of $556$ hours of interviews, narrations, and talks spanning a broad spectrum of speakers and a rich vocabulary.
@@
Table 1: We train and evaluate on multiple datasets consisting of $720$ hours of gesture clips comprising $7000+$ speakers. For evaluation, we curate task-specific benchmarks from the publicly available AVSpeech [18] dataset.
Dataset split #\# hours #\# spk. avg. clip #\# videos
@@
AVS-Ret test 0.31 404 2.27 500
AVS-Spot test 0.38 384 2.76 500
AVS-Asd test 0.44 398 3.15 500
@@ 4.1 Cross-modal Retrieval
4.2 Gesture Word Spotting
@@ 4.3 Active Speaker Detection
[…]
@@ 4.4 AVSpeech Test Benchmarks
AVS-Spot. To quantitatively evaluate the gesture spotting task, we manually curate a new test dataset where we search and annotate clips that clearly contain a word that is gestured. We obtain $500$ such clips, each containing a target word that is clearly gestured. The manual annotation process removes all kinds of label noise in the test set, allowing for a faithful evaluation of our newly defined gesture spotting task. Additionally, we also manually annotate these target words with binary “stress/emphasis” labels, which can have important cues about the gesture (Table 5). We provide additional annotations of the AVS-Spot test set and more results with it in the supplementary.
AVS-Asd. To build the evaluation dataset for active speaker detection, we first choose $500$ “target” clips. For these target clips, we create three evaluation subsets, where we choose P1P-1 clips from different speakers, where $P=2,4,6$. We report the accuracy of detecting the correct target speaker out of the $P$ different speakers.
@@ 5.1 Baselines
For baselines, we report performance of zero-shot pre-trained vision-language models [29, 64] and pre-trained GestSync [23], which learns gesture-audio correspondences by solving for audio-visual synchronization. We also report scores after fine-tuning all these models further on our training data for a fair comparison. In addition, we compare with the semantic encoder of GestureDiffuCLIP [8], by training it on our dataset.
@@ 5.2 Cross-modal Retrieval
In Table 2, we compare the performance of JEGAL against other baselines on the cross-modal retrieval task. Zero-shot evaluation of foundational vision-language models like LanguageBind [64] and Clip4Clip [29] leads to higher than chance performance. These models are designed to capture different kinds of features: they cannot handle a large number of frames, and learn non-gesture attributes like identity and scene. Fine-tuning these models improves their performance on the task, but it is still far from the performance of JEGAL . GestSync [23] clearly performs better than the foundational vision-language models post-finetuning. However, since this network is trained to detect synchronization offsets in speech and video, its representations perform poorly for global semantic tasks like retrieval. This is also partly true for our model when we turn off the global phrase loss (Table 6 row $2$ vs row $3$). GestureDiffuCLIP’s semantic encoder [8] performs only second best among the baselines. The lack of local word-level semantic supervision leads to an inferior performance compared to JEGAL for both GestSync and GestureDiffuCLIP.
Furthermore, none of the baseline approaches ingest and fuse multi-modal speech-text inputs. Our JEGAL model outperforms previous methods by a large margin. JEGAL can retrieve gestures from speech or text queries with similar performance. The opposite direction is also true, i.e. retrieving speech or text for a query gesture clip. Finally, retrieving with the fused speech-text representation is clearly better than the unimodal variants, showing that the speech and text embeddings each encode information that is not present in the other modality.
@@
Clip4Clip [29] 9.20
Language-Bind [64] 9.40
GestSync [23] 15.63
@@
Clip4Clip [29] 17.59
Language-Bind [64] 15.80
GestSync [23] 21.04
GestureDiffuClip [8] 19.50
@@
JEGAL 61.00
JEGAL 41.80
JEGAL 63.60
5.3 Gesture Word Spotting
@@ 5.4 Active Speaker Detection
In Table 4, we show the accuracy of identifying the target speaker for a given text and/or speech segment. This task is different from our other tasks – it does not need a strong holistic understanding of the gesture sequence like retrieval, nor does it need semantic word-level understanding. It can simply be solved by checking for frame-level synchronization, which is exactly why we see GestSync [23] performs the best on this task. None of the other models are trained with strong frame-level video-speech supervision, and hence, perform worse. JEGAL comes at a close second after GestSync. We also see that speech information is more useful here compared to text.
@@
Clip4Clip [29] 62.6 39.6 31.8
Language-Bind [64] 51.4 32.2 23.8
GestSync [23] 54.2 31.6 23.8
@@
Clip4Clip [29] 63.8 42.3 32.8
Language-Bind [64] 56.8 38.7 30.1
GestSync [23] 81.2 64.8 54.4
GestureDiffuClip [8] 61.8 39.4 28.6
@@
JEGAL 65.6 44.4 34.6
JEGAL 74.4 50.0 40.4
JEGAL 76.8 57.8 48.0
@@ 6.1 Speech v/s Text Modalities
We have already seen how our model can flexibly leverage speech or text modality to solve three different kinds of tasks. In Fig 4, we show an example where audio-based gesture word spotting is successful, but text-based spotting is not. We see that the uttered word “action” has been emphasized (using pitch graph). This example leads us to do a deeper analysis of gesture word spotting based on stress cues. We divide the binary stress labels to split the AVS-Spot test set ($500$ samples) into two subsets, one containing only samples with stressed/emphasized words ($100$ samples) and the other subset containing the remaining words ($400$ samples). In Table 5, we report the same spotting accuracy metric on these subsets separately. While we saw previously in Table 3 that speech-based gesture spotting is clearly inferior compared to text-based gesture spotting, it is not always the case. We see that the difference in spotting accuracy between the stressed and non-stressed words is higher for the speech modality. This indicates that speech modality pays more attention to emphasis of the word compared to the text modality.
Table 5: Stressed words are more likely to be spotted with speech-based spotting than non-stressed words. As seen in the column $\Delta$”, the difference between stressed vs. non-stressed word spotting is higher for the speech modality.